A sample of a 355mL drink has a molarity of 0.17M. Calculate the amount of sugar (sucrose) it contains and identify the drink.
1
Expert's answer
2020-06-15T14:34:19-0400

m(C12H22O11) = 342.3 × 0.17 = 58.2 g

C = 58.2 / 0.355 = 164 g/L

Mountain Dew